OpenStudy (anonymous):

I need help with my Application of Differentiation homework, Anyone want to help me? Please do!! :(

OpenStudy (anonymous):

xlnx-2x x>0

OpenStudy (atlas):

whats the question?

OpenStudy (anonymous):

xlnx-2x x>0

OpenStudy (anonymous):

oh you have to find the coordinates of the min and max points

OpenStudy (atlas):

is it xlnx -2x ; x>0

OpenStudy (atlas):

then differentiate the given expression. At maxima and minima the df/dx =0

OpenStudy (anonymous):

yeah but how do I differentiate xlnx

OpenStudy (atlas):

oh don't you know how to differentiate product functions: d(f(x)g(x))/dx = f(x)g'(x) + g(x)f'(x)

OpenStudy (atlas):

x d(lnx)/dx + lnx dx/dx

OpenStudy (atlas):

To find dlnx/dx; Assume lnx =y ; So dy/dx is what u want.......right? or i can write x = e^y or dx/dx = d(e^y)/dx or 1 = {d(e^y)/dy}*dy/dx

OpenStudy (atlas):

I guess u can take the work up frm here :P

OpenStudy (anonymous):

so.. do I differentiate xlnx and -2x seperately and add them or multiply them.. ? :

OpenStudy (anonymous):

.differentiation confuses me and so that makes my homework tonight more difficult, but I don't have time to go back over differentiation cause I find it hard and I've a load to do.

OpenStudy (atlas):

when differentiation acts on sum of functions it is quite simple: d(f(x)+g(x))/dx =f'(x) + g'(x)

OpenStudy (atlas):

d(xlnx -2x)/dx = d(xlnx)/dx -d(2x)/dx

OpenStudy (atlas):

@yaysocks : don't worry you will get used to it.......everyone becomes comfortable after a point

OpenStudy (anonymous):

I'm sorry, I know I'm completely clueless, but now I have lnx=1, how do I get the turning point(s)?

OpenStudy (anonymous):

Thank you for your patience!

OpenStudy (anonymous):

and help of course!